OpenStudy (anonymous):

The radian measure of an angle that is 105 degrees 32' 45'' is? a. 1.59 radians b. 1.84 radians c. 0.59 radians d. 2.78 radians

OpenStudy (unklerhaukus):

\[2\pi^r=360°\]\[\pi^r=180°\] \[\frac{\pi^r}{180°}=1\] \[\theta°=\theta°\times1=\theta°\times\frac{\pi^r}{180°}=\frac{\theta}{180}\times\pi^r\]

OpenStudy (unklerhaukus):

\[105°32'45''=105°+32'\times\frac{1°}{60'}+45''\times\frac{1°}{60'}\times\frac{1'}{60''}\]